OpenStudy (emac):

suppose xy=-1 and dy/dt=4. Find dx/dt when x=-3

OpenStudy (drummergirl7):

can u help me plz

OpenStudy (sshayer):

xy=-1 \[x \frac{ dy }{ dt }+y \frac{ dx }{ dt }=0\] when x=-3 -3*y=-1 \[y=\frac{ -1 }{ -3 }=\frac{ 1 }{ 3 }\] \[\frac{ dy }{ dt }=4\] plug all the values i.e.,x=-3 y=1/3 dy/dt=1/3 then find dx/dt

OpenStudy (sshayer):

or you can do like this \[x \frac{ dy }{ dt }+y \frac{ dx }{ dt }=0\] multiply by x \[x^2\frac{ dy }{ dt }+xy \frac{ dx }{ dt }=0\] \[(-3)^2*4+(-1)\frac{ dx }{ dt }=0,\] \[\frac{ dx }{ dt }=?\]

OpenStudy (emac):

Okay that makes sense! I got 36 which is the correct answer. Thank you!!!